OpenStudy (anonymous):

what is the discriminant of 6x^2-5x=0

Parth (parthkohli):

\[\rm{discriminant = (coefficient~ of ~x)^2 - 4\times(coefficient ~of~ x^2) \times (coefficient ~of~x^0) }\]

Parth (parthkohli):

In other words,\[D = b^2 - 4ac\]where\[ax^2 + bx + c = 0\]is your equation.

Parth (parthkohli):

The value of \(a\) is \(6\), \(b\) is \(-5\) and \(c = 0\). Continue?

OpenStudy (anonymous):

in simplify form

Parth (parthkohli):

Eh? You have to find \(5^2 - 4\times6\times0\)

OpenStudy (anonymous):

6x^2-5x=0

Parth (parthkohli):

Yeah, but the discriminant of that is \(5^2 - (4\times6\times0)\). Can you simplify that?

OpenStudy (anonymous):

yes i need simplify form of that

Parth (parthkohli):

Yeah, so you simplify it.